### v1

# Semantic Integrity Constraints: Declarative Guardrails for AI-Augmented Data Processing Systems

Abstract.
The emergence of AI-augmented Data Processing Systems (DPSs) has introduced powerful semantic operators that extend traditional data management capabilities with LLM-based processing. However, these systems face fundamental reliability (a.k.a. trust) challenges, as LLMs can generate erroneous outputs, limiting their adoption in critical domains. Existing approaches to LLM constraints—ranging from user-defined functions to constrained decoding—are fragmented, imperative, and lack semantics-aware integration into query execution. To address this gap, we introduce Semantic Integrity Constraints (SICs), a novel declarative abstraction that extends traditional database integrity constraints to govern and optimize semantic operators within DPSs. SICs integrate seamlessly into the relational model, allowing users to specify common classes of constraints (e.g., grounding and soundness) while enabling query-aware enforcement and optimization strategies.
In this paper, we present the core design of SICs, describe their formal integration into query execution, and detail our conception of grounding constraints, a key SIC class that ensures factual consistency of generated outputs. In addition, we explore novel enforcement mechanisms, combining proactive (constrained decoding) and reactive (validation and recovery) techniques to optimize efficiency and reliability. Our work establishes SICs as a foundational framework for trustworthy, high-performance AI-augmented data processing, paving the way for future research in constraint-driven optimizations, adaptive enforcement, and enterprise-scale deployments.
1. Introduction
Large Language Models (LLMs) have transformed the field of data management in the last couple of years, extending traditional systems with semantic processing capabilities over structured and unstructured data. In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Shankar et al., 2024a; Liu et al., 2025; Anderson et al., 2025; Lu et al., 2025). We broadly refer to these systems as AI-augmented Data Processing Systems (DPSs). While AI-augmented DPSs introduce novel ways to process data, their use of LLMs lead to potential reliability issues that manifest into phenomena like hallucinations. These reliability issues hinder the adoption of AI-augmented DPSs in high-stakes sectors, such as healthcare, law, and finance. For these systems to be fully adopted, guardrails must be placed around semantic operators.
LLM constraints—also known as assertions or guardrails—have been actively explored in academia and industry (Beurer-Kellner et al., 2023; Singhvi et al., 2024; Rebedea et al., 2023; Shankar et al., 2024a, b; Dong et al., 2024a; Liu et al., 2024; Shankar et al., 2024c; AI, 2023a, b; Index, 2023). These constraints leverage user-defined functions (UDFs) or LLM validators to check an LLM’s output, evaluating code or natural language predicates, respectively. A related area of work involves constrained decoding, where constraints are applied to the LLM during its generation phase to ensure the output matches a specified structure (Beurer-Kellner et al., 2023; Dong et al., 2024b; guidance, 2022; OpenAI, 2024b; Willard and Louf, 2023; Zheng et al., 2024). While these existing approaches have proven their utility in enhancing the reliability of LLM outputs, they are imperative, fragmented, and lack awareness of operator semantics, limiting their usability and efficiency.
To address these limitations, we introduce Semantic Integrity Constraints (SICs) as a unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs. Unlike traditional integrity constraints that prevent invalid modifications to the database, SICs protect against erroneous outputs from semantic operators. Similar to traditional integrity constraints, SICs make the specification of constraints for predominant use cases easy and more advanced cases possible. They are declarative, allowing for a rich set of enforcement and optimization techniques, such as constraint pushdown, constrained decoding, fused evaluation, and adaptive retries. In addition, SICs are expressed in the same language that defines data processing and integrate seamlessly into the relational model. These design decisions are in line with previous efforts that argue for statement and solution specifications on LLM-based system components (Stoica et al., 2024) as well as an (extended) relational model in AI-augmented DPSs (Liu et al., 2025; Patel et al., 2024; Madden et al., 2024; Lu et al., 2025).
The design of the right interface for SICs and the optimization of their enforcement present many exciting research opportunities and challenges. Different declarative interfaces have been proposed for semantic queries, with the majority adopting the relational model (Liu et al., 2025; Patel et al., 2024; Lu et al., 2025). Although the community has not yet converged on an exact query language, we believe that the language for declaring semantic queries and SICs must be developed in tandem to maintain consistency and tight integration. Moreover, the constraint interface must enable users to declare SICs for the majority of use cases in a simple manner that does not require a deep understanding of the underlying implementation details. Towards this goal, the community will need to conduct user surveys to ensure that design decisions are grounded in real application requirements and do not introduce leaky abstractions. Lastly, optimizing SIC enforcement will require integrating modern techniques in LLM assertions and constrained decoding with traditional database query optimization methods. To meet ever-growing performance demands, novel approaches and extensions to existing solutions must also be developed. These solutions have the potential to accelerate the adoption of AI-augmented solutions in enterprise applications.
This paper presents our vision for SICs and sketches a general SIC framework for relational data processing with unstructured data. Section 2 introduces our proposed constraint interface, followed by Section 3, a categorization of SICs that addresses the most common use cases for LLM constraints. In Section 4, we discuss the enforcement of SICs, detail various optimizations, and highlight multiple promising research directions. Section 5 focuses on grounding constraints, a class of SICs that we are implementing in VectraFlow (Lu et al., 2025), an AI-augmented data flow system developed at Brown University. Finally, we outline several interesting open problems in Section 6.
2. Constraint Interface
We propose an interface that adopts the relational model and enables users to define SICs with the same declarative language used for semantic queries. This approach aligns with traditional methods for expressing integrity constraints, where users declare the logic of a constraint without detailing its physical implementation, allowing the DPS to choose the most efficient way to enforce SICs.
The majority of AI-augmented DPSs provide data flow interfaces, where users chain operators to define a query (Liu et al., 2025; Patel et al., 2024; Anderson et al., 2025). These systems offer a pandas-like interface in Python that exposes semantic operators separately from standard operators. This logical separation prevents declaring both semantic and standard predicates or functions within the same operator. However, recent work has introduced piped data flow syntax to SQL, enhancing its flexibility and extensibility (Shute et al., 2024). This new SQL dialect provides a basis for a unified interface between semantic and traditional queries, while also offering integrity constraints as first class citizens. Even though we design this interface for relational AI-augmented DPSs (Liu et al., 2025; Patel et al., 2024; Lu et al., 2025), we expect that it can extend to DPSs with other data models too (Shankar et al., 2024a; Anderson et al., 2025).
LABEL:lst:query shows an example query that incorporates semantic functionality into pipe SQL syntax. We first describe the high-level query and then discuss the SIC-specific code later in this and the following sections. The query is declared over the ehr_table (line 1), a relation of electronic health records (EHRs). Each tuple in the relation corresponds to a patient, and we assume that it contains attributes for the patient’s ehr, biological sex, and age. The query first canonicalizes the sex attribute’s value into “Male” or “Female” (line 2) because the value could have been in another representation, e.g., “M” or “F”. The SET expression is defined using a string prefixed with the letter “p”. This expression declares a prompt string literal (p-string), which specifies functions or predicates that the DPS evaluates with an LLM. Similar to Python’s formatted string literals (f-strings), p-strings can refer to other attributes in the query within curly braces. Lines 5 to 7 extract the patient’s physical exam, lab results, and medical history from the ehr, and line 8 summarizes the patient’s medical history. Because the EXTEND operator effectively invokes a semantic map to project new attributes, each projected attribute declares a name and type using the AS keyword. Lastly, the query filters tuples by selecting patients likely to have sepsis (lines 9 to 10). Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the data flow (line 15). Further note that the order of operations specified in the query does not determine the execution order. For example, the DPS should execute the filter on line 9 before summarizing the patient’s medical history on line 8 to avoid excessive computation.
Semantic query with pipe SQL syntax and SICs.
Crucially, our constraint interface extends the ASSERT operator provided by pipe SQL syntax, allowing users to declare SICs anywhere in the data flow. In the example query, assert operators are used in lines 4 and 11 and onward. These operators specify predicates that can refer to all attributes and operators defined upstream. This design enables the assert operator to validate values produced by attribute-generating semantic operators (e.g., maps, aggregates) as well as the presence or absence of tuples from nonattribute-generating operators (e.g., filters, joins, top-).
Since the predicates of the declared assert operators may not be satisfied, the DPS can retry semantic operators in the case of failure. For each semantic operator, the user can specify a retry threshold, which is the maximum number of times the DPS can retry an operator with the same input. The user can also set a retry threshold that applies to all semantic operators in the data flow by default. In LABEL:lst:query, the retry threshold for the SET operator is defined in line 3, and all other operators use the default threshold.
The user can also specify a failure mode (CONTINUE, DROP, or ABORT) for each assert operator, which determines how the DPS handles the failed input tuple(s). Similar to the retry threshold, the user can set a default failure mode. In the example query, line 4 defines the CONTINUE failure mode on the assert operator, while all other assert operators use the default. If a tuple does not satisfy the assert operator’s predicate, then the CONTINUE failure mode continues to propagate it downstream, while DROP drops it. Note that CONTINUE and DROP correspond to the (soft) Suggest and (hard) Assert constructs in DSPy Assertions (Singhvi et al., 2024), respectively. Lastly, ABORT aborts the entire query if a tuple does not satisfy the predicate. While the semantics of aborting are straightforward in a transactional context, in a data flow context, this means that the system will not only drop the tuple that violated the SIC but also discard all of its dependent tuples, such as those that would be in the same aggregation window. In any case, the DPS will log the appropriate errors and warnings when a tuple fails a SIC or when tuples are dropped to notify users of failures.
3. Constraint Classes
We categorize SICs into different classes (Table 1) that are designed to capture the predominant use cases for LLM constraints:
| Class | Description | Examples |
| Domain | Value is from a specified domain | Data types, structured output, multiple choices, numerical ranges, or length constraints |
| Inclusion/Exclusion | Value incl/excl terms, content, tones, styles, etc. | Inclusion of attending physician names or exclusion of PII from EHRs |
| Grounding | Value is grounded in the source values | Extracted results or summaries are factual w.r.t. original EHRs |
| Soundness | LLM’s reasoning process is logically sound | Predictions of patient diagnoses are based on sound reasoning |
| Relevance | Value is relevant to the given task | Extracted histories from EHRs are medical (desired) rather than social (undesired) |
| Assertions | Arbitrary predicate holds in the data flow | All included URLs in generated research reports lead to valid websites |
Domain Constraints. In the context of SICs, domain constraints remain similar to their traditional counterpart, checking whether an attribute’s value is from a specified domain. For example, a domain constraint can enforce that a value is of a certain data type or structure, within a numerical range, from a finite set of possible values, or less than a certain length. (Type definitions of generated attributes from semantic operators implicitly specify domain constraints.) In the example query from LABEL:lst:query, the first conjunct in line 4 and the conjunct in line 14 explicitly define domain constraints, restricting the sex to either “Male” or “Female” and the summary’s length to less than 1000 characters, respectively. This class of SICs comprise the majority of constraint use cases (Liu et al., 2024). Importantly, the DPS can enforce domain constraints programmatically without relying on an LLM validator, providing correctness guarantees on their evaluation.
Inclusion and Exclusion (IE) Constraints. IE constraints determine if particular terms, content, topics, or abstract features like tones and styles are included or excluded from an attribute’s value. For instance, these constraints ensure that an attending physician’s name is present in an EHR, while avoiding the inclusion of PII. Users can specify these constraints with a set of values that indicate what must be included or excluded. This set can be derived from a fixed collection or another attribute’s value. The user can opt for checks based on exact keyword matches or semantic similarity, even using regular expressions for matches on emails or phone numbers. Users can also define IE constraints with a p-string for content, topics, tones, or styles, as shown in line 13 of the example query, which excludes test results from the medical history summary. In this case, the DPS employs an LLM to enforce the constraint, while lightweight models can be used to handle PII or toxic content checks, such as NER models, LSTMs, or feature-based classifiers.
Grounding Constraints. These constraints target both extractive and abstractive use cases, where users want to check whether the output values of semantic operators are grounded in the source values. For instance, line 11 from the example query ensures that the extracted physical exam and lab results appear verbatim in the original EHR (extractive use case). Similarly, line 12 validates that the generated summary of the patient’s medical history is factually consistent with the EHR (abstractive use case). We discuss grounding constraints in greater detail in Section 5.
Soundness Constraints. These constraints verify whether the underlying LLM’s reasoning process is logically sound in that its premises are true and its reasoning steps are valid. This class of constraints targets nonattribute-generating semantic operators and semantic transformations from one or more values to a new value. For such cases, it is nontrivial to simply evaluate predicates on the operator’s output without invoking an LLM validator to redo the task. Instead of this similarly error-prone approach, the DPS uses an LLM validator to check the underlying LLM’s reasoning process, e.g., chain-of-thought (CoT) (Wei et al., 2022). For example, consider the semantic filter (defined on lines 9 and 10), which selects tuples where the patient is predicted to have sepsis based on their medical condition. Assume that the query earlier retrieved relevant sepsis-related medical information (omitted for brevity). This retrieved information is considered to outline sepsis symptoms and the diagnostic rationale. Under the hood, the DPS prompts the filter’s LLM to return its CoT in addition to a boolean value that represents the selection result. The CoT first lists the premises, i.e., the patient’s medical condition and the retrieved medical information relevant to sepsis. The premises are followed by the reasoning steps that lead to the selection result. Then, the input and output of the filter’s LLM are given to an LLM validator to ensure that the premises are grounded in the input and that the reasoning steps are valid. This step is defined with the assert operator in line 15. Similarly, a soundness constraint is specified in the second conjunct of line 4, which ensures that sex is canonicalized correctly. If the underlying LLM already performs well in a closed-book setting and the retrieved documents do not provide enough context to produce a valid answer, another approach is to have the LLM generate its CoT first and then verify each proposed premise with the input tuple(s) or documents retrieved afterwards. The DPS can also use more sophisticated CoT prompting and verification techniques, which leverage structured natural language formats or programming languages (Ling et al., 2023; Lyu et al., 2023; Chen et al., 2023). This is an avenue for future research.
Relevance Constraints. These constraints validate if an attribute’s value is relevant for the given semantic operator’s task; i.e., they verify whether the underlying LLM followed the instructions prompted by the user. Since relevance constraints are always appropriate, the DPS can enforce them on all generated attributes by default. We assume that this is the case in the example query; e.g., the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than their social history. Relevance constraints are evaluated by passing the operator’s prompt, input, and output to an LLM validator for verification.
Assertions. Like assertions in traditional integrity constraints, these constraints enable users to specify arbitrary predicates that must hold in the data flow. The DPS cannot offer many optimizations for assertions due to their black-box nature; however, such constraints give users the flexibility to express conditions that are not supported by the primitives described previously.
4. Constraint Enforcement
Once users define SICs, the DPS will identify the most efficient enforcement method. The DPS initially moves all specified SICs to the earliest evaluation point to minimize unnecessary computation, akin to predicate pushdown in logical query optimization. If a SIC involves checking attributes, the DPS will move the constraint to the earliest data flow point where all attributes are defined, usually right after a semantic map or aggregate operator, since these are the only attribute-generating operators in our extended relational model. For SICs that verify operator processing logic, like soundness constraints, the DPS positions them at their associated semantic operator. Additionally, the DPS converts specified SICs into text and appends them to the LLM’s prompt, ensuring that the semantic operator’s LLM is aware of the constraints on its results. At this point, the DPS will follow either a proactive or reactive constraint enforcement strategy.
4.1. Proactive Enforcement
The DPS can proactively enforce certain types of SICs by pushing their enforcement into the LLM’s decoding process via constrained decoding (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022). In doing so, the output of the LLM will be guaranteed to satisfy the specified constraints by construction; thus, requiring no further checks on its output. Provided the constrained decoding overhead remains low, the DPS should proactively enforce SICs where feasible to optimize performance. Concerns around distorting the LLM’s distribution can be addressed with recent methods as well (Park et al., 2024).
Grammar-Based Constrained Decoding. An effective method for using constrained decoding to ensure specific SICs is by encoding them as a formal grammar. This is similar to how libraries (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) allow users to provide a JSON Schema (Org, 2010) to which the LLM’s output adheres. These libraries convert the JSON Schema into a context-free grammar or regular expression and apply a push-down or finite automaton, respectively, to guide the output. This allows the DPS to enforce certain domain constraints. Beyond exploring more efficient grammar-based decoding methods, open challenges include determining which SICs can be expressed as grammars and how to represent and compile multiple SICs into a grammar if they are defined on the same attribute.
Direct Constrained Decoding. An alternative constrained decoding method is to enforce SICs directly, without translating them into a formal grammar. Although we do not anticipate that this will allow the DPS to enforce all SICs, we expect it to enable a broader range of constraints than grammar-based methods. Previous research (Beurer-Kellner et al., 2023) introduced a direct, more flexible approach, but is significantly slower than current grammar-based techniques (Zheng et al., 2024). Whether this more flexible enforcement method can be implemented to efficiently enforce SICs is an open problem.
4.2. Reactive Enforcement and Recovery Actions
The DPS can also reactively enforce SICs by first detecting a constraint violation and then performing a recovery action.
4.2.1. Violation Detection
To detect a SIC violation, the DPS simply evaluates its associated predicate and checks if the predicate returns false. However, if multiple SICs are specified in the data flow, there are many opportunities to optimize their collective evaluation.
Constraint Reordering. If the DPS needs to execute multiple SICs sequentially (e.g., if the user specifies the constraints on the same attribute), then it will reorder the constraints to reduce execution costs. This is similar to predicate reordering in traditional database optimization. For example, the DPS can reorder SICs based on the cost and selectivity of each SIC’s predicate (Abadi et al., 2003). The cost of a SIC can be based on a single metric, such as execution time, or even a linear combination of cost metrics. The selectivity of a SIC is the probability that an output tuple fails the associated predicate. The DPS can learn the cost and selectivity of each SIC via offline or online methods by computing lightweight statistics.
Constraint Fusion. The DPS will further improve performance by merging SICs together into a single execution unit for certain scenarios. Akin to operator fusion, the DPS will combine code-based SICs together to reduce intermediate data materialization and unnecessary memory traffic. The DPS will also do the same for LLM-based SICs, where several constraints are evaluated by a single LLM invocation. Though, as more LLM-based constraints are fused together, the accuracy of results from the LLM validator is likely to degrade due to the increase in task complexity and input length (Levy et al., 2024). As such, finding the optimal number of constraints to combine that maintains target accuracy metrics while minimizing execution costs is an open problem.
4.2.2. Violation Recovery
To recover from a SIC violation, the DPS will retry the semantic operator with the failed input tuple(s) by using prompt or model-based techniques to improve the output quality. We outline a few approaches for retrying semantic operators. These approaches can be combined in different ways, and it is interesting to explore whether the DPS can learn their optimal combination with respect to different target metrics, such as accuracy, latency, and monetary cost.
Few-Shot Prompting. A common approach for retrying a semantic operator is to append the erroneous output and corresponding feedback message to the previous prompt on each failure (Singhvi et al., 2024; Shankar et al., 2024a). The feedback message can be a static error message specified for a code-based predicate or part of the response from an LLM validator. This refinement process provides the underlying LLM negative few-shot examples of the given task for in-context learning (Brown et al., 2020).
A complementary approach is to use the top- most similar positive few-shot examples, which are input-output pairs that previously satisfied all the specified SICs for the semantic operator. Importantly, examples are retrieved on the basis of their similarity to the LLM’s current input. The DPS can use a cache with a vector index to efficiently store and retrieve these examples. For each retry, the number of retrieved examples can increase to improve the quality of the LLM’s output. We believe that this approach will be particularly useful in data flow systems, where queries are mostly static. There are several challenges in our setting, e.g., determining ’s initial value and increment size for each retry, and example cache management policies.
Chain-of-Thought Prompting. When retrying a semantic operator, the DPS can also prompt the underlying LLM to use more reasoning steps in its CoT. By effectively scaling up inference-time computation, recent work has shown that long CoT in reasoning models generally improves the accuracy of responses (OpenAI, 2024a; DeepSeek-AI, 2025). For each retry, the DPS can thus prompt the LLM to use increasingly longer CoT to generate better outputs. It is unclear how long the initial CoT should be and by how much to increase it on each retry.
Model-Based Improvements. To improve the results of semantic operators, the DPS can invoke a more powerful LLM to generate a more accurate output, or use multiple LLMs to adopt a majority voting approach, e.g., where the same input is sent to increasingly more LLMs on each retry. In simple cases, such as when the operator outputs a classification label or a numerical value, the DPS can programmatically consider the outputs of each LLM and return the most common result or their average, respectively. In more complex scenarios, such as when the operator generates a summary, the DPS can use a separate LLM to consider all the outputs and produce the final result. An open avenue for exploration is to compare the accuracy of these two methods when various cost metrics are fixed.
5. Grounding Constraints
In this section, we describe grounding constraints in more detail to illustrate the design decisions required to integrate SICs. These constraints are being implemented in VectraFlow (Lu et al., 2025), an AI-augmented data flow system that we are developing at Brown. However, we aim to keep the presentation general since the ideas proposed here are applicable to other AI-augmented DPSs as well. We first discuss preliminaries related to grounding constraints and then describe their implementation in the rest of the section.
To understand the rationale for grounding constraints, we note a key distinction: an LLM’s response comes from its internal (parametric) knowledge, external (nonparametric) knowledge, or a combination of both. In an AI-augmented DPS, this external knowledge includes input tuples to the system or tuples retrieved from stored tables, termed source tuples, with their elements called source values. These values, considered factual, are essential to grounding constraints. Across extractive and abstractive contexts, users seek to confirm if a value generated by a semantic operator is anchored in source values. Grounding constraints facilitate these verifications.
5.1. Enforcement
When a grounding constraint is specified for an attribute, the DPS recursively applies checks to all ancestor attributes in its lineage. These grounding checks are necessary because the attribute may have been derived through a series of intermediate semantic operators from the source value(s). The DPS evaluates these checks immediately after attributes are generated, ensuring that the operator’s output is grounded in its input. In the example query from LABEL:lst:query, med_hist_sum is derived from med_hist, which, in turn, is derived from ehr. Since the grounding constraint specified on line 12 for med_hist_sum recursively defines a grounding check on med_hist, the query does not need to explicitly specify a grounding constraint on med_hist, though doing so is equivalent.
The notion of grounding differs depending on whether the operator is extractive or abstractive. For extractive operators, grounded output is precisely contained within the input. For abstractive operators, grounded output is factually consistent with the input. The DPS cannot inherently identify whether an operator is extractive or abstractive; hence, users must annotate extractive semantic maps. Without such annotations, attribute-generating operators in the data flow are presumed abstractive if they produce an attribute within the grounding constraint’s lineage. In the example query, we see the EXTRACTIVE keyword included for the extractive operations on lines 5 to 7, while no annotation is present for the abstractive operation that summarizes the medical history on line 8. Evaluating grounding checks on extractive and abstractive semantic operators necessitates different check types. The DPS will automatically assess the appropriate checks based on each operator’s semantics. We will now discuss these grounding check types.
5.2. Types of Grounding Checks
We propose three types of grounding checks with different verification mechanisms, targeting either extractive or abstractive tasks.
Exact Match. A grounding check based on exact match verifies whether an output attribute’s value is exactly contained within the input value. These checks are designed for extractive semantic map operations, where users require exact guarantees on output values, such as phys_exam, lab_res, and med_hist in the example query. The DPS can further optimize exact match verification by building inverted indices over stored relations. However, in data flow systems, evaluating these checks on values derived from input data streams are not amenable to such optimizations.
Factual Consistency Score. These checks evaluate whether an output value is grounded in the input value(s) by computing factual consistency scores with a specialized model. Each sentence in an output value is separately given with the input value(s) to the factual consistency model, and the resulting score determines whether the sentence is grounded. While these models are stochastic, they can validate summaries from abstractive semantic maps (e.g., med_hist_sum) and aggregates. Models like MiniCheck have also been shown to be more efficient than their LLM counterparts and reach GPT-4 accuracy (Tang et al., 2024).
LLM Validator. LLM-based grounding checks also target abstractive use cases, invoking an LLM validator to determine grounding. The LLM validator receives the output value and input value(s), and it returns a factual consistency score or a boolean value. Additionally, the LLM validator provides a feedback message, which can help the semantic operator output more grounded responses when retried. In fact, the DPS can combine the LLM-based check with the score-based check, where it invokes the LLM validator only if the factual consistency model classifies the output as ungrounded.
6. Open Problems
In this section, we present additional open problems that pave the way for promising research opportunities.
Adaptive Enforcement. A DPS can strategically manage performance and risk by dynamically applying SICs. By tracking SIC violations for each semantic operator, it determines when checks are necessary, thus reducing overhead. Violations incur a (normalized) penalty score reflecting their risk level, and these scores are combined into a sum. If this sum exceeds a set threshold, a retry is initiated to maintain critical integrity. This approach allows the system to adapt on the basis of each operator’s reliability, concentrating on critical validations while minimizing disruptions when operator confidence is high. Another approach involves adaptively enforcing SICs based on target metrics (e.g., accuracy, F1 score).
LLM Input Verification. While our discussion of SICs has mainly emphasized validating semantic operator outputs, it is equally crucial to validate the inputs, which may come from external sources (including a stored table). This step is necessary to avoid security risks, such as (i) prompt attacks (AI, 2024; Llama, 2024; Microsoft, 2024; Rebedea et al., 2023) or (ii) prevent unauthorized access to dark data (IBM, 2023), which refers to unprotected data hidden in a vector database that should not be used because it can pose security, privacy, or compliance risks due to its sensitive contents. An open challenge here is developing robust and scalable mechanisms to perform real-time, context-aware input verification that can effectively identify and filter malicious or unauthorized data before it reaches the LLM.
Enterprise-Wide Automated Constraint Verification. To enable a “check all” feature for semantic queries, each semantic operator in the query path must be automatically verified. Strategically determining what checks to use, where and how to place them involves examining data flow semantics and lineage to pinpoint where operator outputs affect subsequent results. The types of checks, which range from basic range checks to LLM validators, should match each operator’s semantics and risk of its output. Although our discussion has focused on single queries, enterprise settings often involve numerous queries that share common domain-specific constraints. Therefore, an important challenge is developing a mechanism to automatically capture and manage these recurring constraints in a constraint store and apply them consistently across all relevant semantic queries, ensuring enterprise-wide data integrity. To further enhance this process, the DPS should log all SIC violations, coupled with user feedback on incorrect validation results. This approach facilitates continuous LLM refinement through training, fine-tuning, or in-context learning, ultimately bolstering the accuracy of enterprise-wide constraint verification.
Leveraging LLM Progress. As LLM reliability and efficiency advance, they will influence SIC design and related trade-offs. Increases in LLM accuracy allow for more adaptivity in SIC enforcement. A critical question is: How do we determine when LLM accuracies warrant this shift? Additionally, as LLMs become more efficient, using advanced ensemble methods (e.g., confidence-based aggregation) to further improve reliability as well as maintaining LLM diversity to prevent shared errors become important.
7. Related Work
Database integrity constraints were first described in Codd’s relational model (Codd, 1970) and later implemented in INGRES (Held et al., 1975; Stonebraker, 1975) and System R (Astrahan et al., 1976; Chamberlin et al., 1976; Eswaran and Chamberlin, 1975). Since then, constraints have been supported by all major databases to ensure that modifications maintain data consistency. SICs build on top of traditional integrity constraints, improving the reliability of the semantic operators themselves by protecting against erroneous outputs from the underlying LLMs.
In ML, assertions were introduced as an abstraction to monitor and improve models, invoking UDFs over model outputs to flag possible errors (Kang et al., 2020). Recent works adopt similar ideas by further leveraging ML models and LLM validators to define guardrails on LLMs (Rebedea et al., 2023; Dong et al., 2024a; AI, 2023a, b; Index, 2023), with DSPy Assertions introducing an iterative refinement approach to improve output quality (Singhvi et al., 2024). Methods for validating an LLM’s CoT have been developed as well (Ling et al., 2023). In addition to these reactive constraint enforcement approaches, a related area of work uses constrained decoding to proactively enforce constraints (OpenAI, 2024b; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022; Beurer-Kellner et al., 2023). SICs consolidate these various approaches into a unified and declarative abstraction designed for AI-augmented DPSs. Prior works like spade (Shankar et al., 2024b), which automatically synthesizes assertions, and EvalGen (Shankar et al., 2024c), which aligns LLM validators with human preferences, are complimentary to SICs and can be integrated to enhance usability and reliability.
Several recent systems have augmented traditional database operators with the semantic capabilities of LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025; Lu et al., 2025; Shankar et al., 2024a). So far, DocETL (Shankar et al., 2024a) is the only AI-augmented DPS to provide validation as a first class citizen. They enable users to express simple Python statements and validation prompts for iteratively refining operator outputs. We further their approach by proposing a declarative, relational interface for the easy specification of common constraint classes and their optimized enforcement.
Many methods have been developed to evaluate the factual consistency of LLM responses with respect to provided source documents. This problem is also known as document-level textual entailment. In addition to an exact keyword match approach that is only appropriate for extractive use cases, recent methods for evaluating factual consistency in abstractive tasks are largely categorized as those that use LLM validators and those that use specialized factual consistency models. The latter models are specifically trained to evaluate factual consistency, primarily in the context of summarization (Tang et al., 2024; Zha et al., 2023; Kryscinski et al., 2020; Goyal and Durrett, 2020; Laban et al., 2022). These models provide factual consistency scores and are more efficient than LLM-based methods, with some even reaching GPT-4 accuracy (Tang et al., 2024; Zha et al., 2023). Regardless of the specific method for evaluating factual consistency, the interface provided by SICs is designed to be agnostic of the underlying implementation.
8. Conclusions
SICs are a unified abstraction that extend traditional database integrity constraints to enhance the reliability of AI-augmented DPSs. They provide a declarative, relational interface to easily specify common classes of constraint while maintaining the flexibility to support more sophisticated use cases. Crucially, this declarative interface enables the DPS to apply a wide range of highly optimized implementations. By adopting this principled approach, SICs open new, grounded research avenues towards the goal of leveraging AI-augmented DPSs for critical workloads where reliability is paramount.
Acknowledgements.
We thank Shu Chen, Duo Lu, Malte Schwarzkopf, Weili Shi, and the VectraFlow team for their valuable feedback and contributions.References
- (1)
- Abadi et al. (2003) Daniel J. Abadi, Don Carney, Uğur Çetintemel, Mitch Cherniack, Christian Convey, Sangdon Lee, Michael Stonebraker, Nesime Tatbul, and Stan Zdonik. 2003. Aurora: a new model and architecture for data stream management. The VLDB Journal 12 (2003), 120–139. https://doi.org/10.1007/s00778-003-0095-z
- AI (2023a) Guardrails AI. 2023a. Guardrails. https://github.com/guardrails-ai/guardrails
- AI (2024) Guardrails AI. 2024. Detect Jailbreak. https://hub.guardrailsai.com/validator/guardrails/detect_jailbreak
- AI (2023b) LangChain AI. 2023b. LangChain. https://github.com/langchain-ai/langchain
- Anderson et al. (2025) Eric Anderson, Jonathan Fritz, Austin Lee, Bohou Li, Mark Lindblad, Henry Lindeman, Alex Meyer, Parth Parmar, Tanvi Ranade, Mehul A. Shah, et al. 2025. The Design of an LLM-powered Unstructured Analytics System. In CIDR.
- Astrahan et al. (1976) M. M. Astrahan, M. W. Blasgen, D. D. Chamberlin, K. P. Eswaran, J. N. Gray, P. P. Griffiths, W. F. King, R. A. Lorie, P. R. McJones, J. W. Mehl, G. R. Putzolu, I. L. Traiger, B. W. Wade, and V. Watson. 1976. System R: relational approach to database management. 1, 2 (June 1976), 97–137. https://doi.org/10.1145/320455.320457
- Beurer-Kellner et al. (2023) Luca Beurer-Kellner, Marc Fischer, and Martin Vechev. 2023. Prompting Is Programming: A Query Language for Large Language Models. Proc. ACM Program. Lang. 7, PLDI, Article 186 (June 2023), 24 pages. https://doi.org/10.1145/3591300
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language models are few-shot learners. In Proceedings of the 34th International Conference on Neural Information Processing Systems (Vancouver, BC, Canada) (NIPS ’20). Curran Associates Inc., Red Hook, NY, USA, Article 159, 25 pages.
- Chamberlin et al. (1976) D. D. Chamberlin, M. M. Astrahan, K. P. Eswaran, P. P. Griffiths, R. A. Lorie, J. W. Mehl, P. Reisner, and B. W. Wade. 1976. SEQUEL 2: A Unified Approach to Data Definition, Manipulation, and Control. IBM Journal of Research and Development 20, 6 (1976), 560–575. https://doi.org/10.1147/rd.206.0560
- Chen et al. (2023) Wenhu Chen, Xueguang Ma, Xinyi Wang, and William W. Cohen. 2023. Program of Thoughts Prompting: Disentangling Computation from Reasoning for Numerical Reasoning Tasks. Transactions on Machine Learning Research (2023).
- Codd (1970) E. F. Codd. 1970. A relational model of data for large shared data banks. Commun. ACM 13, 6 (June 1970), 377–387. https://doi.org/10.1145/362384.362685
- DeepSeek-AI (2025) DeepSeek-AI. 2025. DeepSeek-R1: Incentivizing Reasoning Capability in LLMs via Reinforcement Learning. arXiv:2501.12948 [cs.CL] https://arxiv.org/abs/2501.12948
- Dong et al. (2024a) Yi Dong, Ronghui Mu, Gaojie Jin, Yi Qi, Jinwei Hu, Xingyu Zhao, Jie Meng, Wenjie Ruan, and Xiaowei Huang. 2024a. Building Guardrails for Large Language Models. arXiv:2402.01822 [cs.CL] https://arxiv.org/abs/2402.01822
- Dong et al. (2024b) Yixin Dong, Charlie F. Ruan, Yaxing Cai, Ruihang Lai, Ziyi Xu, Yilong Zhao, and Tianqi Chen. 2024b. XGrammar: Flexible and Efficient Structured Generation Engine for Large Language Models. arXiv:2411.15100 [cs.CL] https://arxiv.org/abs/2411.15100
- Eswaran and Chamberlin (1975) Kapali P. Eswaran and Donald D. Chamberlin. 1975. Functional specifications of a subsystem for data base integrity. In Proceedings of the 1st International Conference on Very Large Data Bases (Framingham, Massachusetts) (VLDB ’75). Association for Computing Machinery, New York, NY, USA, 48–68. https://doi.org/10.1145/1282480.1282484
- Goyal and Durrett (2020) Tanya Goyal and Greg Durrett. 2020. Evaluating Factuality in Generation with Dependency-level Entailment. In Findings of the Association for Computational Linguistics: EMNLP 2020, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, Online, 3592–3603. https://doi.org/10.18653/v1/2020.findings-emnlp.322
- guidance (2022) guidance. 2022. guidance-ai. https://github.com/guidance-ai/guidance
- Held et al. (1975) G. D. Held, M. R. Stonebraker, and E. Wong. 1975. INGRES: a relational data base system. In Proceedings of the May 19-22, 1975, National Computer Conference and Exposition (Anaheim, California) (AFIPS ’75). Association for Computing Machinery, New York, NY, USA, 409–416. https://doi.org/10.1145/1499949.1500029
- IBM (2023) IBM. 2023. What is dark data? Retrieved 2025-02-27 from https://www.ibm.com/think/topics/dark-data
- Index (2023) Llama Index. 2023. Llama Index. https://github.com/run-llama/llama_index
- Kang et al. (2020) Daniel Kang, Deepti Raghavan, Peter Bailis, and Matei Zaharia. 2020. Model Assertions for Monitoring and Improving ML Models. In Proceedings of Machine Learning and Systems, Vol. 2. 481–496.
- Kryscinski et al. (2020) Wojciech Kryscinski, Bryan McCann, Caiming Xiong, and Richard Socher. 2020. Evaluating the Factual Consistency of Abstractive Text Summarization. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, Online, 9332–9346. https://doi.org/10.18653/v1/2020.emnlp-main.750
- Laban et al. (2022) Philippe Laban, Tobias Schnabel, Paul N. Bennett, and Marti A. Hearst. 2022. SummaC: Re-Visiting NLI-based Models for Inconsistency Detection in Summarization. Transactions of the Association for Computational Linguistics 10 (2022), 163–177. https://doi.org/10.1162/tacl_a_00453
- Levy et al. (2024) Mosh Levy, Alon Jacoby, and Yoav Goldberg. 2024. Same Task, More Tokens: the Impact of Input Length on the Reasoning Performance of Large Language Models. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). Association for Computational Linguistics, Bangkok, Thailand, 15339–15353. https://doi.org/10.18653/v1/2024.acl-long.818
- Ling et al. (2023) Zhan Ling, Yunhao Fang, Xuanlin Li, Zhiao Huang, Mingu Lee, Roland Memisevic, and Hao Su. 2023. Deductive Verification of Chain-of-Thought Reasoning. In Advances in Neural Information Processing Systems, A. Oh, T. Naumann, A. Globerson, K. Saenko, M. Hardt, and S. Levine (Eds.), Vol. 36. Curran Associates, Inc., 36407–36433. https://proceedings.neurips.cc/paper_files/paper/2023/file/72393bd47a35f5b3bee4c609e7bba733-Paper-Conference.pdf
- Liu et al. (2025) Chunwei Liu, Matthew Russo, Michael Cafarella, Lei Cao, Peter Baille Chen, Zui Chen, Michael Franklin, Tim Kraska, Samuel Madden, Rana Shahout, and Gerardo Vitagliano. 2025. Palimpzest: Optimizing AI-Powered Analytics with Declarative Query Processing. In CIDR.
- Liu et al. (2024) Michael Xieyang Liu, Frederick Liu, Alexander J. Fiannaca, Terry Koo, Lucas Dixon, Michael Terry, and Carrie J. Cai. 2024. ”We Need Structured Output”: Towards User-centered Constraints on Large Language Model Output. In Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI EA ’24). Association for Computing Machinery, New York, NY, USA, Article 10, 9 pages. https://doi.org/10.1145/3613905.3650756
- Llama (2024) Meta Llama. 2024. Prompt Guard. https://www.llama.com/docs/model-cards-and-prompt-formats/prompt-guard/
- Lu et al. (2025) Duo Lu, Siming Feng, Jonathan Zhou, Franco Solleza, Malte Schwarzkopf, and Uğur Çetintemel. 2025. VectraFlow: Integrating Vectors into Stream Processing. In CIDR.
- Lyu et al. (2023) Qing Lyu, Shreya Havaldar, Adam Stein, Li Zhang, Delip Rao, Eric Wong, Marianna Apidianaki, and Chris Callison-Burch. 2023. Faithful Chain-of-Thought Reasoning. arXiv:2301.13379 [cs.CL] https://arxiv.org/abs/2301.13379
- Madden et al. (2024) Samuel Madden, Michael Cafarella, Michael Franklin, and Tim Kraska. 2024. Databases Unbound: Querying All of the World’s Bytes with AI. Proc. VLDB Endow. 17, 12 (Aug. 2024), 4546–4554. https://doi.org/10.14778/3685800.3685916
- Microsoft (2024) Microsoft. 2024. Prompt Shields. https://learn.microsoft.com/en-us/azure/ai-services/content-safety/concepts/jailbreak-detection
- OpenAI (2024a) OpenAI. 2024a. Learning to Reason with LLMs. https://openai.com/index/learning-to-reason-with-llms/
- OpenAI (2024b) OpenAI. 2024b. Structured Outputs. https://platform.openai.com/docs/guides/structured-outputs/
- Org (2010) JSON Schema Org. 2010. JSON Schema Spec. https://github.com/json-schema-org/json-schema-spec
- Park et al. (2024) Kanghee Park, Jiayu Wang, Taylor Berg-Kirkpatrick, Nadia Polikarpova, and Loris D'Antoni. 2024. Grammar-Aligned Decoding. In Advances in Neural Information Processing Systems, A. Globerson, L. Mackey, D. Belgrave, A. Fan, U. Paquet, J. Tomczak, and C. Zhang (Eds.), Vol. 37. Curran Associates, Inc., 24547–24568. https://proceedings.neurips.cc/paper_files/paper/2024/file/2bdc2267c3d7d01523e2e17ac0a754f3-Paper-Conference.pdf
- Patel et al. (2024) Liana Patel, Siddharth Jha, Parth Asawa, Melissa Pan, Carlos Guestrin, and Matei Zaharia. 2024. Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data. arXiv:2407.11418 [cs.DB] https://arxiv.org/abs/2407.11418
- Rebedea et al. (2023) Traian Rebedea, Razvan Dinu, Makesh Narsimhan Sreedhar, Christopher Parisien, and Jonathan Cohen. 2023. NeMo Guardrails: A Toolkit for Controllable and Safe LLM Applications with Programmable Rails. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Yansong Feng and Els Lefever (Eds.). Association for Computational Linguistics, Singapore, 431–445. https://doi.org/10.18653/v1/2023.emnlp-demo.40
- Shankar et al. (2024a) Shreya Shankar, Tristan Chambers, Tarak Shah, Aditya G. Parameswaran, and Eugene Wu. 2024a. DocETL: Agentic Query Rewriting and Evaluation for Complex Document Processing. arXiv:2410.12189 [cs.DB] https://arxiv.org/abs/2410.12189
- Shankar et al. (2024b) Shreya Shankar, Haotian Li, Parth Asawa, Madelon Hulsebos, Yiming Lin, J. D. Zamfirescu-Pereira, Harrison Chase, Will Fu-Hinthorn, Aditya G. Parameswaran, and Eugene Wu. 2024b. spade: Synthesizing Data Quality Assertions for Large Language Model Pipelines. Proc. VLDB Endow. 17, 12 (Aug. 2024), 4173–4186. https://doi.org/10.14778/3685800.3685835
- Shankar et al. (2024c) Shreya Shankar, J.D. Zamfirescu-Pereira, Bjoern Hartmann, Aditya Parameswaran, and Ian Arawjo. 2024c. Who Validates the Validators? Aligning LLM-Assisted Evaluation of LLM Outputs with Human Preferences. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 131, 14 pages. https://doi.org/10.1145/3654777.3676450
- Shute et al. (2024) Jeff Shute, Shannon Bales, Matthew Brown, Jean-Daniel Browne, Brandon Dolphin, Romit Kudtarkar, Andrey Litvinov, Jingchi Ma, John Morcos, Michael Shen, David Wilhite, Xi Wu, and Lulan Yu. 2024. SQL Has Problems. We Can Fix Them: Pipe Syntax In SQL. Proc. VLDB Endow. 17, 12 (Aug. 2024), 4051–4063. https://doi.org/10.14778/3685800.3685826
- Singhvi et al. (2024) Arnav Singhvi, Manish Shetty, Shangyin Tan, Christopher Potts, Koushik Sen, Matei Zaharia, and Omar Khattab. 2024. DSPy Assertions: Computational Constraints for Self-Refining Language Model Pipelines. arXiv:2312.13382 [cs.CL] https://arxiv.org/abs/2312.13382
- Stoica et al. (2024) Ion Stoica, Matei Zaharia, Joseph Gonzalez, Ken Goldberg, Koushik Sen, Hao Zhang, Anastasios Angelopoulos, Shishir G. Patil, Lingjiao Chen, Wei-Lin Chiang, and Jared Q. Davis. 2024. Specifications: The missing link to making the development of LLM systems an engineering discipline. arXiv:2412.05299 [cs.SE] https://arxiv.org/abs/2412.05299
- Stonebraker (1975) Michael Stonebraker. 1975. Implementation of integrity constraints and views by query modification. In Proceedings of the 1975 ACM SIGMOD International Conference on Management of Data (San Jose, California) (SIGMOD ’75). Association for Computing Machinery, New York, NY, USA, 65–78. https://doi.org/10.1145/500080.500091
- Tang et al. (2024) Liyan Tang, Philippe Laban, and Greg Durrett. 2024. MiniCheck: Efficient Fact-Checking of LLMs on Grounding Documents. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, Yaser Al-Onaizan, Mohit Bansal, and Yun-Nung Chen (Eds.). Association for Computational Linguistics, Miami, Florida, USA, 8818–8847. https://doi.org/10.18653/v1/2024.emnlp-main.499
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, brian ichter, Fei Xia, Ed Chi, Quoc V Le, and Denny Zhou. 2022. Chain-of-Thought Prompting Elicits Reasoning in Large Language Models. In Advances in Neural Information Processing Systems, S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Eds.), Vol. 35. Curran Associates, Inc., 24824–24837. https://proceedings.neurips.cc/paper_files/paper/2022/file/9d5609613524ecf4f15af0f7b31abca4-Paper-Conference.pdf
- Willard and Louf (2023) Brandon T. Willard and Rémi Louf. 2023. Efficient Guided Generation for Large Language Models. arXiv:2307.09702 [cs.CL] https://arxiv.org/abs/2307.09702
- Zha et al. (2023) Yuheng Zha, Yichi Yang, Ruichen Li, and Zhiting Hu. 2023. AlignScore: Evaluating Factual Consistency with A Unified Alignment Function. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 11328–11348. https://doi.org/10.18653/v1/2023.acl-long.634
- Zheng et al. (2024) Lianmin Zheng, Liangsheng Yin, Zhiqiang Xie, Chuyue Sun, Jeff Huang, Cody Hao Yu, Shiyi Cao, Christos Kozyrakis, Ion Stoica, Joseph E. Gonzalez, Clark Barrett, and Ying Sheng. 2024. SGLang: Efficient Execution of Structured Language Model Programs. arXiv:2312.07104 [cs.AI] https://arxiv.org/abs/2312.07104