### v2

# Semantic Integrity Constraints: Declarative Guardrails for AI-Augmented Data Processing Systems

Abstract.
AI-augmented data processing systems (DPSs) integrate large language models (LLMs) into query pipelines, allowing powerful semantic operations on structured and unstructured data. However, the reliability (a.k.a. trust) of these systems is fundamentally challenged by the potential for LLMs to produce errors, limiting their adoption in critical domains. To help address this reliability bottleneck, we introduce semantic integrity constraints (SICs)—a declarative abstraction for specifying and enforcing correctness conditions over LLM outputs in semantic queries. SICs generalize traditional database integrity constraints to semantic settings, supporting common types of constraints, such as grounding, soundness, and exclusion, with both proactive and reactive enforcement strategies.
We argue that SICs provide a foundation for building reliable and auditable AI-augmented data systems. Specifically, we present a system design for integrating SICs into query planning and runtime execution and discuss its realization in AI-augmented DPSs. To guide and evaluate the vision, we outline several design goals—covering criteria around expressiveness, runtime semantics, integration, performance, and enterprise-scale applicability—and discuss how our framework addresses each, along with open research challenges.
1. Introduction
Large language models (LLMs) have transformed the field of data management in the last couple of years, extending traditional systems with semantic processing capabilities over structured and unstructured data. In particular, several projects have integrated semantic operators into database and data flow systems by augmenting traditional operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025). We broadly refer to these systems as AI-augmented data processing systems (DPSs). While AI-augmented DPSs introduce novel ways to process data, their use of LLMs lead to potential reliability issues that manifest into phenomena like hallucinations. These reliability issues hinder the adoption of AI-augmented DPSs in high-stakes sectors, such as healthcare, law, and finance. Without guardrails, it will be difficult to fully adopt these systems.
LLM constraints—also known as assertions or guardrails—have recently been actively explored in academia and industry (Beurer-Kellner et al., 2023; Singhvi et al., 2024; Rebedea et al., 2023; Shankar et al., 2024a, b; Dong et al., 2024a; Liu et al., 2024; Shankar et al., 2024c; AI, 2023a, b; Index, 2023; TruLens, 2021; Zheng et al., 2023). These constraints leverage user-defined functions (UDFs) or LLM validators to check an LLM’s output, evaluating code or natural language predicates, respectively. A related approach is constrained decoding, where constraints are applied during the LLM’s generation to ensure that the output adheres to a specified structure (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023). While these approaches have proven their utility in enhancing LLM reliability, they are imperative and fragmented, limiting their usability.
We propose semantic integrity constraints (SICs) as a declarative, unified abstraction that extends traditional database integrity constraints to support AI-augmented DPSs. Unlike traditional integrity constraints that prevent invalid modifications to the database, SICs protect against erroneous outputs from semantic operators. Similar to their traditional counterparts, SICs make defining constraints for predominant use cases easy and more advanced cases possible, aligning with recent work that argues for specifications on LLM-based system components (Stoica et al., 2024).
| Success Criterion | Description |
| Expressiveness | SICs support a range of constraint types—including domain, grounding, and soundness—through a declarative interface (Section 3) |
| Executable Semantics | Each constraint has precise execution semantics—including what is checked, when it is enforced, and what actions are taken on failure—ensuring semantically consistent system behavior and predictable outcomes for users (Section 4) |
| System Integration | SICs integrate with query interfaces, optimizers, and execution engines in AI-augmented DPSs (Section 2) |
| Performance Awareness | Enforcement strategy is selected based on cost and reliability across proactive or reactive methods (Section 4.2) |
| Enterprise-Scale Applicability | SICs support observability, reuse, and auditability across workflows, with mechanisms for conflict detection and policy-driven enforcement at organizational scale (Sections 5 and 6) |
This paper presents our vision for SICs and sketches a general framework for integrating them into AI-augmented DPSs. To evaluate the feasibility and impact of this vision, we define a set of design goals and success criteria: (1) expressiveness, to capture a broad range of common constraint types; (2) executable semantics, with well-defined runtime behavior; (3) system integration, via seamless support within data flow and query processing pipelines; (4) performance-aware enforcement, balancing reliability and cost; and (5) enterprise-scale applicability, including observability, constraint reuse, conflict detection, and auditability. These criteria, outlined in Table 1, were defined through our partnership with the Rhode Island Hospital. They guide our design and are reflected in the structure of the paper. Section 3.1 introduces the SIC interface, while Section 3.2 categorizes SICs across key enforcement classes aligned with real-world LLM use cases. Section 4 explores enforcement strategies, spanning logical and physical planning, and outlines cost-aware mechanisms. Section 5 discusses observability components. Finally, Section 6 presents open challenges in large-scale constraint management.
2. System Model
We begin with an overview of the system architecture (Figure 1) for supporting SICs and defer a more detailed discussion of the core components to later sections. First, the user expresses their semantic query along with SICs via the query interface (Section 3.1). This interface forwards the query’s text to the query parser, which produces a corresponding logical query plan. Next, the logical optimizer (Section 4.1) applies logical rewrite rules to the initial plan, yielding an optimized logical plan. This plan is then passed to the physical optimizer (Section 4.2), which considers different physical implementations of the plan and selects the optimal one. The resulting optimized physical plan is then executed by the execution engine. To enable performance monitoring, an observability stack is integrated with the execution engine (Section 5). Finally, the constraint store recommends and analyzes SICs (Section 6).
System architecture for SICs. Orange circles and turquoise diamonds represent semantic operators and SICs, respectively. Green squares indicate annotations that describe the physical implementation of each node.
3. Constraint Specification
We suggest defining SICs declaratively, allowing users to focus on constraint logic while the optimizer determines the best execution method. We outline the main components of our proposed interface and detail classes of constraints relevant to semantic data processing. User surveys should be conducted to ensure usability.
3.1. Interface
To provide a unified interface for queries and SICs, users declare constraints in the query language. Specifically, we present a relational interface that extends pipe SQL syntax (Shute et al., 2024) with semantic operators (e.g., map, aggregate, filter, join, top-) and SICs. Though, SICs are compatible with any declarative AI-augmented DPS. The interface allows users to specify two types of information to the DPS for each constraint: (1) what the constraint is and (2) how the DPS should react when the constraint is violated.
Declaring Constraints. SICs specify predicates either on generated tuple attributes from semantic operators or on the semantic operators themselves. We make this distinction because while some operators (e.g., map, aggregate) generate attributes, other operators (e.g., filter, join, top-) do not; yet, constraints should still be defined in either case. The syntax for declaring a constraint on an attribute involves a SQL expression starting with ASSERT <attribute>, followed by a keyword that refers to the constraint class (Section 3.2) and any necessary parameters. A constraint on an operator is similar: ASSERT <operator>. Users can also combine constraints via logical connectives (e.g., ASSERT <pred1> AND <pred2>).
Retry Thresholds. Users can define a retry threshold for each operator. When a constraint associated with the operator (either on the operator or on an attribute generated by the operator) is violated, the DPS retries the operator until the output satisfies the constraint or the retry threshold is reached. To specify a retry threshold, users write RETRY <threshold> after the operator.
Failure Modes. Failure modes (CONTINUE, IGNORE, or ABORT) are also associated with operators and specify what the DPS should do when a failing output reaches the retry threshold. With CONTINUE, the DPS continues processing the query, allowing errors to propagate downstream. IGNORE specifies that the DPS should ignore invalid tuples in the processing (including dependent tuples, such as those that would be in the same aggregation window). Lastly, ABORT aborts the entire query. To specify a failure mode, the user writes <mode> ON FAIL after the retry threshold.
End-To-End Example. LABEL:lst:query shows an example semantic query with various SICs. The query is declared over the ehr_table (line 1), a relation of electronic health records (EHRs). Each tuple in the relation corresponds to a patient, and we assume that it contains attributes for the patient’s ehr and dob. The query first canonicalizes the dob attribute’s value into YYYY-MM-DD (line 2) because the value could have been in another format. The expression is defined using a string prefixed with the letter “p”, which declares a prompt string. Prompt strings specify natural language functions or predicates over attributes and are evaluated using LLMs. Line 4 computes the patient’s age using standard SQL functions. Lines 5-7 extract the patient’s physical exam, lab results, and medical history from the ehr, and line 8 summarizes the patient’s medical history. Lastly, the query filters tuples by selecting patients likely to have sepsis (lines 10-11). Here, the AS keyword is also used to alias the filter operator so that it can be referenced later in the query (line 16).
The first constraint is declared on line 3, and the remaining constraints are specified on line 12 and onward. If the summarization step for med_hist_sum (line 8) fails its constraints (lines 13-15), the DPS retries the operator once and then CONTINUEs to propagate the tuple even if it still fails the constraints (line 9). Since no retry thresholds and failure modes are defined in the rest of the query, we assume that some default setting is used. We explain each constraint in more detail in the next section.
Semantic query with pipe SQL syntax and SICs.
3.2. Constraint Classes
We categorize SICs into different classes (Table 2) that are designed to capture the predominant use cases for LLM constraints. We motivate each constraint class using concrete examples from the EHR query (among others) and briefly outline their implementation.
Domain Constraints. In many queries, users want the generated attributes to belong to a specified domain. This includes use cases such as enforcing that a value is of a certain data type or structure, from a set of values, or less than a certain length. For example, one may check whether the predicted prescription dose from a semantic map is within a range of possible values. We extend traditional domain constraints to semantic data processing. In LABEL:lst:query, lines 3 and 15 specify domain constraints, restricting the dob to the regex representation of YYYY-MM-DD and the summary’s length to less than 1000 characters, respectively. The type definitions (e.g., STRING) of generated attributes implicitly define domain constraints. This class of SICs comprises the majority of constraint use cases (Liu et al., 2024) and is enforced deterministically.
Inclusion and Exclusion (IE) Constraints. Users sometimes expect that an attribute generated by an operator includes or excludes certain terms, content, topics, or abstract features such as tones or styles. For example, given a semantic map that summarizes an EHR, users may want to exclude all patient PII from the summary while including the names of attending physicians. IE constraints address this class of predicates. In the example query, line 14 defines an exclusion constraint on med_hist_sum, which excludes test results from the medical history summary. When an IE constraint is defined with a prompt string (as in the example), LLMs are used to check for a semantic match. Users can also specify IE constraints using strings, regular expressions, or even a collection of them (e.g., a set of toxic words to exclude from the output). In such cases, deterministic methods are used to evaluate matches.
Grounding Constraints. Semantic operators, such as maps or aggregates, can extract data or summarize from their inputs. However, since LLMs can hallucinate, users want to ensure that the generated data is grounded in source data. Grounding constraints enable users to enforce such invariants on generated attributes. For example, line 12 ensures that the extracted phys_exam and lab_res are exactly contained in the source EHR. Line 13 asserts that the summarized med_hist_sum is factually consistent with the source EHR. In particular, the last grounding constraint implicitly defines a grounding constraint on med_hist, since med_hist_sum is derived from med_hist, which, in turn, is derived from ehr. Formally, if a grounding constraint is defined on an attribute, the DPS will recursively enforce grounding constraints on all ancestor attributes in its lineage. Thus, each constraint only needs to compare its operator’s input and output to ensure grounding in the source data. A grounding constraint is implemented in one of two ways, depending on whether the specified attribute is generated via an extractive (map) or abstractive (map or aggregate) operator. In the extractive case, the constraint is implemented deterministically and ensures that the operator’s output is precisely contained in its input. In the abstractive case, the constraint is implemented stochastically (e.g., with an LLM or a fact-checking model (Tang et al., 2024)) and ensures that the operator’s summarized output is factually consistent with the input. Users can annotate relevant operators with the EXTRACTIVE (lines 5-7) or ABSTRACTIVE (line 8) keywords to inform the DPS of their semantics; otherwise, the DPS infers these annotations by reasoning about the operator’s prompt string with an LLM.
| Class | Description | Examples |
| Domain | Value is from a specified domain | Data types, structured output, multiple choices, numerical ranges, or length constraints |
| Inclusion/Exclusion | Value incl/excl terms, content, tones, styles, etc. | Inclusion of attending physician names or exclusion of PII from EHR summaries |
| Grounding | Value is grounded in the source values | Extracted test results or summaries are factual w.r.t. original EHRs |
| Soundness | LLM’s reasoning process is logically sound | Predictions of patient diagnoses are based on sound reasoning |
| Relevance | Value is relevant to the given task | Extracted histories from EHRs are medical (desired) rather than social (undesired) |
| Assertions | Arbitrary predicate holds in the data flow | All included URLs in generated research reports lead to valid websites |
Soundness Constraints. To check whether the reasoning of operators themselves is logically sound, users can declare soundness constraints. Although applicable to all operators, this constraint class is particularly useful for operators that do not generate attributes, such as filter, join, or top-. To enforce soundness constraints, the DPS uses LLMs to check the operator’s reasoning process by modifying the operator to output its reasoning steps, e.g., chain-of-thought (CoT) (Wei et al., 2022). Line 16 declares a soundness constraint on the sepsis filter. Assume that the query earlier retrieved relevant sepsis-related medical information (omitted for brevity). This retrieved information is considered to describe the symptoms of sepsis and the diagnostic rationale. Under the hood, the DPS prompts the filter’s LLM to return its CoT in addition to a boolean value that represents the selection result. The CoT first lists the premises, i.e., the patient’s medical condition and the retrieved medical information relevant to sepsis. The premises are followed by the reasoning steps that lead to the selection result. Then, the input and output of the filter’s LLM are given to an LLM validator (or several of them) to ensure that the premises are grounded in the input and that the reasoning steps are valid. The DPS can also use more sophisticated CoT prompting and verification techniques, which use structured natural language formats or programming languages (Ling et al., 2023; Lyu et al., 2023; Chen et al., 2023).
Relevance Constraints. These constraints validate if an attribute’s value is relevant for the given operator’s task; i.e., they verify whether the underlying LLM followed the instructions prompted by the user. Since relevance constraints are always appropriate, the DPS can enforce them on all generated attributes by default. We assume that this is the case in the example query. For example, the DPS checks whether the medical history extracted in line 7 is in fact the patient’s medical history rather than something irrelevant, such as their social history. These constraints can be evaluated by passing the operator’s prompt, input, and output to an LLM judge.
Assertions. Like assertions in traditional integrity constraints, these constraints enable users to specify arbitrary predicates that must hold in the data flow. The DPS cannot offer many optimizations for assertions due to their black-box nature; however, such constraints give users the flexibility to express conditions that are not supported by the primitives described previously.
4. Constraint Enforcement
Constraint enforcement consists of two phases: detection and recovery. First, a constraint detects any violations in the operator’s output. Upon detection of a violation, the operator attempts to recover. This distinction aids in reasoning about optimizing constraint enforcement, as discussed in this section.
4.1. Logical Optimization
The logical optimizer takes as input a logical query plan represented as a data flow graph, where each node is either an operator or a constraint. The goal of the logical optimizer is to minimize the cost of the plan by applying rewrite rules to semantic operators and constraints. Here, we refer to cost as any arbitrary cost metric, such as latency or monetary cost. To keep the presentation focused, we only discuss logical optimizations for constraints.
Constraint Pushdown. First, the optimizer moves all constraints to their earliest evaluation point in the data flow (i.e., immediately following their corresponding operators). This allows the DPS to detect and recover from constraint violations as soon as possible rather than unnecessarily processing erroneous outputs. The optimizer also appends the textual representation of each constraint to their corresponding operator’s LLM prompt, ensuring that the underlying model is aware of the constraints on its results.
Constraint Reordering. The optimizer then reorders the SICs for the same operator. Following prior work (Abadi et al., 2003), the reordering is based on the cost and selectivity of each constraint, where selectivity is defined as the probability that a tuple satisfies the constraint’s predicate. Cost and selectivity estimates can be computed from sampled data. We discuss more details on estimation in Section 4.2.2.
4.2. Physical Optimization
Given an optimized logical plan, the physical optimizer considers various implementations of the plan and returns the optimal one. We begin with a discussion of different approaches to constraint enforcement, and then outline ideas for optimized plan selection.
4.2.1. Enforcement Implementations
Constraints can be enforced in multiple ways, typically categorized as reactive or proactive, based on whether enforcement occurs after or during response generation, respectively. While the DPS can enforce all constraints reactively, only certain ones can be enforced proactively.
Reactive Enforcement. After the LLM produces its entire response, there are different methods to detect a constraint violation. For example, a grounding constraint on an extractive operator can simply check if the operator’s output attribute is contained in the input. On the other hand, a grounding constraint on an abstractive operator can invoke a single LLM judge (Zheng et al., 2023), a fact-checking model (Tang et al., 2024), a compound AI system (Chern et al., 2023), or anything in between. Similarly, there are different approaches for an operator to recover from a violation by retrying. The common approach used in current systems is iterative refinement with feedback (Singhvi et al., 2024; Shankar et al., 2024a; Madaan et al., 2023). Furthermore, the operator can retrieve additional semantically similar few-shot examples (Brown et al., 2020) of prior input-output pairs that satisfied the constraint. These examples can be stored in a cache built on a vector index. Other approaches include retrying with more powerful models, larger ensemble sizes, or higher thinking budgets (Anthropic, 2025; Google, 2025).
Proactive Enforcement. Proactively enforcing constraints can lower costs significantly. For example, constrained decoding (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023) pushes enforcement into the LLM’s decoding process, identifying and masking invalid tokens to ensure compliance by construction. Various libraries (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; guidance, 2022) support JSON Schema (Org, 2010) adherence in LLM outputs by converting the schema into a grammar and then an automaton for token validation. However, constrained decoding is not limited to grammar-based approaches. For instance, using a suffix automaton is more efficient than regex to constrain an LLM to output a substring of its input (Koo et al., 2024), enabling proactive enforcement of grounding constraints. Yet, open source libraries currently do not offer this functionality.
An open challenge is to identify which other SICs can be enforced via constrained decoding. Moreover, developing a framework to efficiently apply various constraints on different output attributes also remains unresolved. Prior research (Beurer-Kellner et al., 2023) has seen some advancement but remains slower than modern grammar-based techniques (Zheng et al., 2024). Nonetheless, it provides a foundation for future exploration. Importantly, we suggest that the DPS should limit constrained decoding to simple constraints (seen by the LLM during training) to avoid sampling unlikely token distributions. This rationale also underlies why constraints are included in the LLM’s prompt before decoding.
Besides deterministic constrained decoding, the DPS can proactively enforce stochastic constraints by asynchronously detecting violations in partial streamed outputs. This approach is similar to previous work (Santhanam et al., 2024), except that the operator now needs to recover if its partial output violates a constraint. For instance, the DPS can proactively enforce a grounding constraint on an abstractive operator by asynchronously validating generated summary sentences with a fact-checking model (Tang et al., 2024). Upon a violation, the operator can backtrack to the decoding step right before the violation. Alternative techniques like re-prompting with partial feedback are also worth considering to address this problem.
4.2.2. Optimized Plan Selection
For a fixed set of constraints and their implementations, the physical optimizer can adapt existing techniques (Liu et al., 2025b; Russo et al., 2025) to select an optimal query plan defined by user-specified cost or reliability (a.k.a. quality) thresholds (omitted from the query interface for brevity). Specifically, rather than measuring the reliability of an operator’s implementation in terms of validation data or the best implementation from prior beliefs (e.g., a “champion model”), reliability is instead defined by the proportion of output tuples that satisfy all of the operator’s constraints (after necessary retries). However, the challenge is selecting each constraint’s implementation to begin with.
As discussed in Section 4.2.1, there are many ways to detect a constraint violation, each with different cost-reliability trade-offs. Formally, the cost of a constraint’s implementation is the overhead of detecting a violation in a single output. The reliability of a constraint’s implementation is defined as the precision and recall of detecting a violation. A deterministic constraint is completely reliable (i.e., 100% precision and recall), while a stochastic constraint is not completely reliable, due to invoking models such as LLMs.
To choose the best constraint implementation, we must estimate costs and reliability for each option. While cost estimation is often simple and can be done during runtime with sampled data, it might not always be necessary. For instance, proactively enforcing a deterministic constraint through constrained decoding is usually quicker than doing it reactively.
Assessing the reliability of stochastic constraints is more challenging. Unlike deterministic constraints that are always reliable, the reliability of stochastic constraints depends on (1) how they are implemented, (2) the output tuples from upstream operators in the query plan, and (3) the input tuples to the query. Moreover, obtaining robust reliability estimates for constraint implementations can be difficult, as they often need ground-truth labels from human annotators. We present several strategies to address this issue.
One strategy is to use historical data and workloads to estimate the precision and recall of different implementations for each constraint, providing these estimates to the physical optimizer for use at runtime. This approach neglects the dependencies mentioned earlier, leading to possibly noisy estimates. Since metrics are averaged across diverse query contexts, they do not effectively reflect interactions with upstream operators. Furthermore, data and workload drift may make these estimates obsolete. It is also generally costly for organizations to annotate execution traces.
An alternative approach is to develop a shared foundation model for organizations, where given representations of (1) a constraint’s implementation, (2) the implementation of the query plan up until the constraint, and (3) the underlying data, the model predicts the precision and recall for the constraint’s implementation. Organizations can optionally fine-tune the model for their use cases. Recent work on foundation database models (Wehrstein et al., 2025) are applicable here, with extensions to address the complexities of semantic processing.
A third approach is to use the model’s confidence score as a proxy for constraint reliability, allowing the optimizer to estimate the reliability (i.e., expected confidence) of a constraint’s implementation at runtime based on sampled data. Rather than defining precision and recall thresholds, users specify confidence thresholds instead. Although many conventional classifiers inherently provide confidence scores, they can be derived from LLMs by applying softmax to the “True” and “False” token logits. The difficulty lies in achieving calibrated confidence scores that accurately represent the implementation’s true correctness probability. Studies (Kadavath et al., 2022; Liu et al., 2025a) indicate that methods like temperature scaling (Guo et al., 2017) can enhance calibration in binary classification tasks. Though, further research is needed to obtain highly-calibrated constraints.
5. Observability
After a user submits a query to the DPS, the observability stack enables the user to understand the performance of their query’s operators and constraints. The stack has three components. The observability interface allows the user to view relevant metrics (e.g., cost, reliability) and inspect query results. The observability store maintains metrics and outputs from each stage in the query. Lastly, the labeling interface enables human annotators to provide ground-truth labels to compute the precision and recall of each stochastic constraint. While methods such as LLM-as-a-judge offer a scalable approximation of human preferences (Zheng et al., 2023) and have become widespread in both academia and industry (Shankar et al., 2024b, c, a; Singhvi et al., 2024; Madaan et al., 2023; AI, 2023a; Rebedea et al., 2023; AI, 2023b; Dong et al., 2024a; Index, 2023; TruLens, 2021), their reliability remains uncertain. As a result, labeling interfaces remain essential. A key open challenge lies in ensuring that the observability store can support the proposed interfaces efficiently at scale and that these interfaces are effectively designed. We detail each component of the stack below.
Observability Interface. Given a user-specified query ID, the observability interface immediately shows metrics such as cost, operator reliability, and constraint selectivity. However, to observe the reliability for stochastic constraints, the user must submit a labeling request for their query. Once the labeling is complete, the interface displays the precision and recall of each constraint. To avoid confusion, the interface indicates whether a constraint is deterministic or stochastic; though, the query interface can expose this distinction through the EXPLAIN command as well. The observability interface also allows the user to inspect individual output tuples. A tuple is annotated if any upstream operator violated a constraint but CONTINUEed to propagate the error. The user can view the tuple’s lineage for more fine-grained debugging.
Observability Store. The observability store is modeled as additional relations in the DPS. It maintains the lineage of each query, enabling systematic debugging and performance monitoring. To obtain human annotations, the DPS can draw concepts from crowd-sourced databases (Franklin et al., 2011; Marcus et al., 2011; Park et al., 2012), such as crowd-sourced attributes and user interface generation; however, the DPS need not implement these ideas in their full generality. For example, suppose the DPS stores a relation where each tuple represents an invocation of a constraint. The relation includes attributes for the constraint’s input and output (i.e., predicted label) among others (e.g., query ID). In particular, the relation also has a crowd-sourced attribute for the true label (initially null). When the user initiates a labeling request, the observability interface submits a query to the DPS, which computes (over a sample of the relation) the precision and recall of each constraint for the given query ID. The query populates each sampled true label by sending tasks to a crowd-sourcing platform.
Labeling Interface. A key aspect of crowd-sourced databases is their ability to automatically create user interfaces for crowd workers based on declarative queries. Here, the DPS produces a labeling interface for annotators, which displays a description of the constraint and its inputs. The annotator’s role is to determine if the inputs meet the constraint. To avoid bias, the original predicted label should be hidden.
6. Enterprise-Wide Constraints
The introduction of SICs into an enterprise environment presents both new opportunities and challenges. To effectively manage SICs within an organization, we propose the development of a constraint store and outline its key desired features.
Auto-Guardrails and Recommendations. Although declaring SICs is relatively straightforward, determining the applicable constraints is more difficult. This requires anticipating the potential failure cases for each semantic operator in the query. Recent systems have been developed to assist users in formulating constraints (Shankar et al., 2024b, c), but these approaches typically generate suggestions based solely on the contents of the user’s query. In enterprise settings, constraint recommendations can be greatly enriched. Organizations typically enforce domain-specific rules that necessitate certain constraints for all pertinent queries. For instance, a hospital might require excluding PII in queries unless essential for analysis. Moreover, different teams within the same organization might use similar queries with shared constraints. Thus, the constraint store needs to automatically capture and manage common constraints organization-wide, facilitating their reuse and recommendation in related queries. We suggest using text embeddings (OpenAI, 2022; Reimers and Gurevych, 2019) and graph embeddings (Xu et al., 2019; Kipf and Welling, 2017; Veličković et al., 2018) to represent constraints and queries. This strategy boosts the efficiency of similarity searches (Malkov and Yashunin, 2020) for identifying relevant constraints, which can then be refined by an LLM to match the user’s context.
Conflicts. Conflicts can arise when queries contain multiple SICs (e.g., from accumulating shared constraints across an organization’s workloads). For example, a query that requests a grounded list of the patient’s medical conditions with full details while also requiring the exclusion of PII may encounter conflicts when the relevant clinical information (e.g., diagnosis dates, provider names, hospital locations) are themselves considered PII, making it infeasible to satisfy both constraints without redaction or abstraction. Likewise, domain constraints over disjoint value sets can lead to inevitable failures. The constraint store allows for conflict analysis using static reasoning and LLM-based methods, issuing warnings. Users can address conflicts by adjusting priorities (e.g., making one constraint as soft), relaxing scope, or merging constraints as needed. This approach builds on early research in active databases (Widom, 1996; Hanson, 1996; Dayal et al., 1988), which suggested techniques for detecting rule conflicts and execution issues, such as non-termination, before execution (Aiken et al., 1992).
7. Related Work
Database integrity constraints were first described in Codd’s relational model (Codd, 1970) and later implemented in INGRES (Held et al., 1975; Stonebraker, 1975) and System R (Astrahan et al., 1976; Chamberlin et al., 1976; Eswaran and Chamberlin, 1975). Since then, constraints have been supported by all major databases.
ML assertions were introduced to monitor and enhance model performance by applying UDFs to model outputs to find potential errors (Kang et al., 2020). Building upon this idea, recent developments broaden assertions for LLMs by employing LLMs as judges and ML classifiers for both offline evaluation (TruLens, 2021; Zheng et al., 2023; Shankar et al., 2024b, c; AI, 2023b; Index, 2023) and runtime guardrails (AI, 2023a; Rebedea et al., 2023; Dong et al., 2024a; Singhvi et al., 2024; Liu et al., 2024). Tools like spade (Shankar et al., 2024b) and EvalGen (Shankar et al., 2024c) guide users by suggesting relevant assertions and their implementations (e.g., Python functions, LLM grader prompts). These tools complement SICs and can suggest declarative constraints to boost enforcement reliability and efficiency. Our vision for enterprise constraints also draws inspiration from these efforts. DSPy Assertions enable users to define assertions in LLM pipelines using UDFs, which compile optimized prompts and iteratively refine outputs at runtime (Singhvi et al., 2024). While SICs use iterative refinement, they do not currently include prompt optimization capabilities, although they can be added. To the best of our knowledge, no framework provides a fully declarative method for optimized constraint enforcement.
A related area of work is constrained decoding, which proactively enforces constraints so that the LLM’s response adheres to them by construction (OpenAI, 2024; Dong et al., 2024b; Willard and Louf, 2023; Zheng et al., 2024; Beurer-Kellner et al., 2023; guidance, 2022; Koo et al., 2024; Kuchnik et al., 2023). Current frameworks focus on grammar-based approaches; however, similar to earlier work (Beurer-Kellner et al., 2023), we view constrained decoding as a more general approach that is not limited to expressing constraints as grammars.
Several recent systems have augmented traditional database operators with LLMs (Patel et al., 2024; Anderson et al., 2025; Liu et al., 2025b; Lu et al., 2025; Shankar et al., 2024a; Dai et al., 2024; Wang and Li, 2025; Urban and Binnig, 2024; Russo et al., 2025). So far, DocETL is the only AI-augmented DPS to provide validation as a first class citizen (Shankar et al., 2024a). They enable users to express simple Python statements and validation prompts for iteratively refining operator outputs. However, their approach to validation is largely imperative. Palimpzest does not support constraints, but they allow users to specify quality thresholds to select optimized physical plans (Liu et al., 2025b; Russo et al., 2025).
8. Conclusions
SICs offer a unified, declarative foundation for improving the reliability of AI-augmented data processing. By generalizing traditional integrity constraints to the semantic behaviors of LLM-augmented operators, SICs bridge the gap between declarative query processing and the uncertainty of generative models. Their integration into the planning and execution layers of modern DPSs enables not only reliable enforcement but also cost-aware optimization, system-wide coordination, and organizational reuse. Through this framework, we outline a principled path toward building auditable, trustworthy AI-augmented systems capable of supporting mission-critical workflows. SICs are not only a guardrail mechanism, but also a catalyst for a new class of optimizations, abstractions, and research directions at the intersection of databases and LLMs.
Acknowledgements.
We thank Shu Chen, Duo Lu, Malte Schwarzkopf, Weili Shi, and the VectraFlow team for their valuable feedback and contributions.References
- (1)
- Abadi et al. (2003) Daniel J. Abadi, Don Carney, Uğur Çetintemel, Mitch Cherniack, Christian Convey, Sangdon Lee, Michael Stonebraker, Nesime Tatbul, and Stan Zdonik. 2003. Aurora: a new model and architecture for data stream management. The VLDB Journal 12 (2003), 120–139. https://doi.org/10.1007/s00778-003-0095-z
- AI (2023a) Guardrails AI. 2023a. Guardrails. https://github.com/guardrails-ai/guardrails
- AI (2023b) LangChain AI. 2023b. LangChain. https://github.com/langchain-ai/langchain
- Aiken et al. (1992) Alexander Aiken, Jennifer Widom, and Joseph M Hellerstein. 1992. Behavior of database production rules: Termination, confluence, and observable determinism. In Proceedings of the 1992 ACM SIGMOD international conference on Management of data. 59–68.
- Anderson et al. (2025) Eric Anderson, Jonathan Fritz, Austin Lee, Bohou Li, Mark Lindblad, Henry Lindeman, Alex Meyer, Parth Parmar, Tanvi Ranade, Mehul A. Shah, et al. 2025. The Design of an LLM-powered Unstructured Analytics System. In CIDR.
- Anthropic (2025) Anthropic. 2025. Claude’s extended thinking. https://www.anthropic.com/news/visible-extended-thinking
- Astrahan et al. (1976) M. M. Astrahan, M. W. Blasgen, D. D. Chamberlin, K. P. Eswaran, J. N. Gray, P. P. Griffiths, W. F. King, R. A. Lorie, P. R. McJones, J. W. Mehl, G. R. Putzolu, I. L. Traiger, B. W. Wade, and V. Watson. 1976. System R: relational approach to database management. 1, 2 (June 1976), 97–137. https://doi.org/10.1145/320455.320457
- Beurer-Kellner et al. (2023) Luca Beurer-Kellner, Marc Fischer, and Martin Vechev. 2023. Prompting Is Programming: A Query Language for Large Language Models. Proc. ACM Program. Lang. 7, PLDI, Article 186 (June 2023), 24 pages. https://doi.org/10.1145/3591300
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language models are few-shot learners. In Proceedings of the 34th International Conference on Neural Information Processing Systems (Vancouver, BC, Canada) (NIPS ’20). Curran Associates Inc., Red Hook, NY, USA, Article 159, 25 pages.
- Chamberlin et al. (1976) D. D. Chamberlin, M. M. Astrahan, K. P. Eswaran, P. P. Griffiths, R. A. Lorie, J. W. Mehl, P. Reisner, and B. W. Wade. 1976. SEQUEL 2: A Unified Approach to Data Definition, Manipulation, and Control. IBM Journal of Research and Development 20, 6 (1976), 560–575. https://doi.org/10.1147/rd.206.0560
- Chen et al. (2023) Wenhu Chen, Xueguang Ma, Xinyi Wang, and William W. Cohen. 2023. Program of Thoughts Prompting: Disentangling Computation from Reasoning for Numerical Reasoning Tasks. Transactions on Machine Learning Research (2023).
- Chern et al. (2023) I-Chun Chern, Steffi Chern, Shiqi Chen, Weizhe Yuan, Kehua Feng, Chunting Zhou, Junxian He, Graham Neubig, and Pengfei Liu. 2023. FacTool: Factuality Detection in Generative AI–A Tool Augmented Framework for Multi-Task and Multi-Domain Scenarios. arXiv:2307.13528 [cs.CL] https://arxiv.org/abs/2307.13528
- Codd (1970) E. F. Codd. 1970. A relational model of data for large shared data banks. Commun. ACM 13, 6 (June 1970), 377–387. https://doi.org/10.1145/362384.362685
- Dai et al. (2024) Hanjun Dai, Bethany Yixin Wang, Xingchen Wan, Bo Dai, Sherry Yang, Azade Nova, Pengcheng Yin, Phitchaya Mangpo Phothilimthana, Charles Sutton, and Dale Schuurmans. 2024. UQE: A Query Engine for Unstructured Databases. In Advances in Neural Information Processing Systems, A. Globerson, L. Mackey, D. Belgrave, A. Fan, U. Paquet, J. Tomczak, and C. Zhang (Eds.), Vol. 37. Curran Associates, Inc., 29807–29838. https://proceedings.neurips.cc/paper_files/paper/2024/file/34b3a40ec9752c1ae48fe85fef8fe8dc-Paper-Conference.pdf
- Dayal et al. (1988) Umeshwar Dayal, Barbara Blaustein, Alex Buchmann, Upen Chakravarthy, Meichun Hsu, R Ledin, Dennis McCarthy, Arnon Rosenthal, Sunil Sarin, Michael J. Carey, et al. 1988. The HiPAC project: Combining active databases and timing constraints. ACM Sigmod Record 17, 1 (1988), 51–70.
- Dong et al. (2024a) Yi Dong, Ronghui Mu, Gaojie Jin, Yi Qi, Jinwei Hu, Xingyu Zhao, Jie Meng, Wenjie Ruan, and Xiaowei Huang. 2024a. Building Guardrails for Large Language Models. arXiv:2402.01822 [cs.CL] https://arxiv.org/abs/2402.01822
- Dong et al. (2024b) Yixin Dong, Charlie F. Ruan, Yaxing Cai, Ruihang Lai, Ziyi Xu, Yilong Zhao, and Tianqi Chen. 2024b. XGrammar: Flexible and Efficient Structured Generation Engine for Large Language Models. arXiv:2411.15100 [cs.CL] https://arxiv.org/abs/2411.15100
- Eswaran and Chamberlin (1975) Kapali P. Eswaran and Donald D. Chamberlin. 1975. Functional specifications of a subsystem for data base integrity. In Proceedings of the 1st International Conference on Very Large Data Bases (Framingham, Massachusetts) (VLDB ’75). Association for Computing Machinery, New York, NY, USA, 48–68. https://doi.org/10.1145/1282480.1282484
- Franklin et al. (2011) Michael J. Franklin, Donald Kossmann, Tim Kraska, Sukriti Ramesh, and Reynold Xin. 2011. CrowdDB: answering queries with crowdsourcing. In Proceedings of the 2011 ACM SIGMOD International Conference on Management of Data (Athens, Greece) (SIGMOD ’11). Association for Computing Machinery, New York, NY, USA, 61–72. https://doi.org/10.1145/1989323.1989331
- Google (2025) Google. 2025. Gemini 2.5 brings enhanced reasoning to enterprise use cases. https://cloud.google.com/blog/products/ai-machine-learning/gemini-2-5-pro-flash-on-vertex-ai
- guidance (2022) guidance. 2022. guidance-ai. https://github.com/guidance-ai/guidance
- Guo et al. (2017) Chuan Guo, Geoff Pleiss, Yu Sun, and Kilian Q. Weinberger. 2017. On calibration of modern neural networks. In Proceedings of the 34th International Conference on Machine Learning - Volume 70 (Sydney, NSW, Australia) (ICML’17). JMLR.org, 1321–1330.
- Hanson (1996) Eric N. Hanson. 1996. The design and implementation of the Ariel active database rule system. IEEE Transactions on Knowledge and Data Engineering 8, 1 (1996), 157–172.
- Held et al. (1975) G. D. Held, M. R. Stonebraker, and E. Wong. 1975. INGRES: a relational data base system. In Proceedings of the May 19-22, 1975, National Computer Conference and Exposition (Anaheim, California) (AFIPS ’75). Association for Computing Machinery, New York, NY, USA, 409–416. https://doi.org/10.1145/1499949.1500029
- Index (2023) Llama Index. 2023. Llama Index. https://github.com/run-llama/llama_index
- Kadavath et al. (2022) Saurav Kadavath, Tom Conerly, Amanda Askell, Tom Henighan, Dawn Drain, Ethan Perez, Nicholas Schiefer, Zac Hatfield-Dodds, Nova DasSarma, Eli Tran-Johnson, Scott Johnston, Sheer El-Showk, Andy Jones, Nelson Elhage, Tristan Hume, Anna Chen, Yuntao Bai, Sam Bowman, Stanislav Fort, Deep Ganguli, Danny Hernandez, Josh Jacobson, Jackson Kernion, Shauna Kravec, Liane Lovitt, Kamal Ndousse, Catherine Olsson, Sam Ringer, Dario Amodei, Tom Brown, Jack Clark, Nicholas Joseph, Ben Mann, Sam McCandlish, Chris Olah, and Jared Kaplan. 2022. Language Models (Mostly) Know What They Know. arXiv:2207.05221 [cs.CL] https://arxiv.org/abs/2207.05221
- Kang et al. (2020) Daniel Kang, Deepti Raghavan, Peter Bailis, and Matei Zaharia. 2020. Model Assertions for Monitoring and Improving ML Models. In Proceedings of Machine Learning and Systems, Vol. 2. 481–496.
- Kipf and Welling (2017) Thomas N. Kipf and Max Welling. 2017. Semi-Supervised Classification with Graph Convolutional Networks. In International Conference on Learning Representations. https://openreview.net/forum?id=SJU4ayYgl
- Koo et al. (2024) Terry Koo, Frederick Liu, and Luheng He. 2024. Automata-based constraints for language model decoding. In First Conference on Language Modeling. https://openreview.net/forum?id=BDBdblmyzY
- Kuchnik et al. (2023) Michael Kuchnik, Virginia Smith, and George Amvrosiadis. 2023. Validating Large Language Models with ReLM. In Proceedings of Machine Learning and Systems, D. Song, M. Carbin, and T. Chen (Eds.), Vol. 5. Curan, 457–476. https://proceedings.mlsys.org/paper_files/paper/2023/file/93c7d9da61ccb2a60ac047e92787c3ef-Paper-mlsys2023.pdf
- Ling et al. (2023) Zhan Ling, Yunhao Fang, Xuanlin Li, Zhiao Huang, Mingu Lee, Roland Memisevic, and Hao Su. 2023. Deductive Verification of Chain-of-Thought Reasoning. In Advances in Neural Information Processing Systems, A. Oh, T. Naumann, A. Globerson, K. Saenko, M. Hardt, and S. Levine (Eds.), Vol. 36. Curran Associates, Inc., 36407–36433. https://proceedings.neurips.cc/paper_files/paper/2023/file/72393bd47a35f5b3bee4c609e7bba733-Paper-Conference.pdf
- Liu et al. (2025b) Chunwei Liu, Matthew Russo, Michael Cafarella, Lei Cao, Peter Baille Chen, Zui Chen, Michael Franklin, Tim Kraska, Samuel Madden, Rana Shahout, and Gerardo Vitagliano. 2025b. Palimpzest: Optimizing AI-Powered Analytics with Declarative Query Processing. In CIDR.
- Liu et al. (2025a) Hongfu Liu, Hengguan Huang, Xiangming Gu, Hao Wang, and Ye Wang. 2025a. On Calibration of LLM-based Guard Models for Reliable Content Moderation. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=wUbum0nd9N
- Liu et al. (2024) Michael Xieyang Liu, Frederick Liu, Alexander J. Fiannaca, Terry Koo, Lucas Dixon, Michael Terry, and Carrie J. Cai. 2024. ”We Need Structured Output”: Towards User-centered Constraints on Large Language Model Output. In Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI EA ’24). Association for Computing Machinery, New York, NY, USA, Article 10, 9 pages. https://doi.org/10.1145/3613905.3650756
- Lu et al. (2025) Duo Lu, Siming Feng, Jonathan Zhou, Franco Solleza, Malte Schwarzkopf, and Uğur Çetintemel. 2025. VectraFlow: Integrating Vectors into Stream Processing. In CIDR.
- Lyu et al. (2023) Qing Lyu, Shreya Havaldar, Adam Stein, Li Zhang, Delip Rao, Eric Wong, Marianna Apidianaki, and Chris Callison-Burch. 2023. Faithful Chain-of-Thought Reasoning. arXiv:2301.13379 [cs.CL] https://arxiv.org/abs/2301.13379
- Madaan et al. (2023) Aman Madaan, Niket Tandon, Prakhar Gupta, Skyler Hallinan, Luyu Gao, Sarah Wiegreffe, Uri Alon, Nouha Dziri, Shrimai Prabhumoye, Yiming Yang, Shashank Gupta, Bodhisattwa Prasad Majumder, Katherine Hermann, Sean Welleck, Amir Yazdanbakhsh, and Peter Clark. 2023. SELF-REFINE: iterative refinement with self-feedback. In Proceedings of the 37th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’23). Curran Associates Inc., Red Hook, NY, USA, Article 2019, 61 pages.
- Malkov and Yashunin (2020) Yu A. Malkov and D. A. Yashunin. 2020. Efficient and Robust Approximate Nearest Neighbor Search Using Hierarchical Navigable Small World Graphs. IEEE Transactions on Pattern Analysis and Machine Intelligence 42, 4 (2020), 824–836. https://doi.org/10.1109/TPAMI.2018.2889473
- Marcus et al. (2011) Adam Marcus, Eugene Wu, David R. Karger, Samuel Madden, and Robert C. Miller. 2011. Crowdsourced Databases: Query Processing with People. In CIDR.
- OpenAI (2022) OpenAI. 2022. Introducing text and code embeddings. https://openai.com/index/introducing-text-and-code-embeddings/
- OpenAI (2024) OpenAI. 2024. Structured Outputs. https://platform.openai.com/docs/guides/structured-outputs/
- Org (2010) JSON Schema Org. 2010. JSON Schema Spec. https://github.com/json-schema-org/json-schema-spec
- Park et al. (2012) Hyunjung Park, Hector Garcia-Molina, Richard Pang, Neoklis Polyzotis, Aditya Parameswaran, and Jennifer Widom. 2012. Deco: a system for declarative crowdsourcing. Proc. VLDB Endow. 5, 12 (Aug. 2012), 1990–1993. https://doi.org/10.14778/2367502.2367555
- Patel et al. (2024) Liana Patel, Siddharth Jha, Parth Asawa, Melissa Pan, Carlos Guestrin, and Matei Zaharia. 2024. Semantic Operators: A Declarative Model for Rich, AI-based Analytics Over Text Data. arXiv:2407.11418 [cs.DB] https://arxiv.org/abs/2407.11418
- Rebedea et al. (2023) Traian Rebedea, Razvan Dinu, Makesh Narsimhan Sreedhar, Christopher Parisien, and Jonathan Cohen. 2023. NeMo Guardrails: A Toolkit for Controllable and Safe LLM Applications with Programmable Rails. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Yansong Feng and Els Lefever (Eds.). Association for Computational Linguistics, Singapore, 431–445. https://doi.org/10.18653/v1/2023.emnlp-demo.40
- Reimers and Gurevych (2019) Nils Reimers and Iryna Gurevych. 2019. Sentence-BERT: Sentence Embeddings using Siamese BERT-Networks. arXiv:1908.10084 [cs.CL] https://arxiv.org/abs/1908.10084
- Russo et al. (2025) Matthew Russo, Sivaprasad Sudhir, Gerardo Vitagliano, Chunwei Liu, Tim Kraska, Samuel Madden, and Michael Cafarella. 2025. Abacus: A Cost-Based Optimizer for Semantic Operator Systems. arXiv:2505.14661 [cs.DB] https://arxiv.org/abs/2505.14661
- Santhanam et al. (2024) Keshav Santhanam, Deepti Raghavan, Muhammad Shahir Rahman, Thejas Venkatesh, Neha Kunjal, Pratiksha Thaker, Philip Levis, and Matei Zaharia. 2024. ALTO: An Efficient Network Orchestrator for Compound AI Systems. In Proceedings of the 4th Workshop on Machine Learning and Systems (Athens, Greece) (EuroMLSys ’24). Association for Computing Machinery, New York, NY, USA, 117–125. https://doi.org/10.1145/3642970.3655844
- Shankar et al. (2024a) Shreya Shankar, Tristan Chambers, Tarak Shah, Aditya G. Parameswaran, and Eugene Wu. 2024a. DocETL: Agentic Query Rewriting and Evaluation for Complex Document Processing. arXiv:2410.12189 [cs.DB] https://arxiv.org/abs/2410.12189
- Shankar et al. (2024b) Shreya Shankar, Haotian Li, Parth Asawa, Madelon Hulsebos, Yiming Lin, J. D. Zamfirescu-Pereira, Harrison Chase, Will Fu-Hinthorn, Aditya G. Parameswaran, and Eugene Wu. 2024b. spade: Synthesizing Data Quality Assertions for Large Language Model Pipelines. Proc. VLDB Endow. 17, 12 (Aug. 2024), 4173–4186. https://doi.org/10.14778/3685800.3685835
- Shankar et al. (2024c) Shreya Shankar, J.D. Zamfirescu-Pereira, Bjoern Hartmann, Aditya Parameswaran, and Ian Arawjo. 2024c. Who Validates the Validators? Aligning LLM-Assisted Evaluation of LLM Outputs with Human Preferences. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 131, 14 pages. https://doi.org/10.1145/3654777.3676450
- Shute et al. (2024) Jeff Shute, Shannon Bales, Matthew Brown, Jean-Daniel Browne, Brandon Dolphin, Romit Kudtarkar, Andrey Litvinov, Jingchi Ma, John Morcos, Michael Shen, David Wilhite, Xi Wu, and Lulan Yu. 2024. SQL Has Problems. We Can Fix Them: Pipe Syntax In SQL. Proc. VLDB Endow. 17, 12 (Aug. 2024), 4051–4063. https://doi.org/10.14778/3685800.3685826
- Singhvi et al. (2024) Arnav Singhvi, Manish Shetty, Shangyin Tan, Christopher Potts, Koushik Sen, Matei Zaharia, and Omar Khattab. 2024. DSPy Assertions: Computational Constraints for Self-Refining Language Model Pipelines. arXiv:2312.13382 [cs.CL] https://arxiv.org/abs/2312.13382
- Stoica et al. (2024) Ion Stoica, Matei Zaharia, Joseph Gonzalez, Ken Goldberg, Koushik Sen, Hao Zhang, Anastasios Angelopoulos, Shishir G. Patil, Lingjiao Chen, Wei-Lin Chiang, and Jared Q. Davis. 2024. Specifications: The missing link to making the development of LLM systems an engineering discipline. arXiv:2412.05299 [cs.SE] https://arxiv.org/abs/2412.05299
- Stonebraker (1975) Michael Stonebraker. 1975. Implementation of integrity constraints and views by query modification. In Proceedings of the 1975 ACM SIGMOD International Conference on Management of Data (San Jose, California) (SIGMOD ’75). Association for Computing Machinery, New York, NY, USA, 65–78. https://doi.org/10.1145/500080.500091
- Tang et al. (2024) Liyan Tang, Philippe Laban, and Greg Durrett. 2024. MiniCheck: Efficient Fact-Checking of LLMs on Grounding Documents. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, Yaser Al-Onaizan, Mohit Bansal, and Yun-Nung Chen (Eds.). Association for Computational Linguistics, Miami, Florida, USA, 8818–8847. https://doi.org/10.18653/v1/2024.emnlp-main.499
- TruLens (2021) TruLens. 2021. TruLens. https://www.trulens.org/
- Urban and Binnig (2024) Matthias Urban and Carsten Binnig. 2024. CAESURA: Language Models as Multi-Model Query Planners. In CIDR.
- Veličković et al. (2018) Petar Veličković, Guillem Cucurull, Arantxa Casanova, Adriana Romero, Pietro Liò, and Yoshua Bengio. 2018. Graph Attention Networks. In International Conference on Learning Representations. https://openreview.net/forum?id=rJXMpikCZ
- Wang and Li (2025) Jiayi Wang and Guoliang Li. 2025. AOP: Automated and Interactive LLM Pipeline Orchestration for Answering Complex Queries. In CIDR.
- Wehrstein et al. (2025) Johannes Wehrstein, Carsten Binnig, Fatma Özcan, Shobha Vasudevan, Yu Gan, and Yawen Wang. 2025. Towards Foundation Database Models. In CIDR.
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H. Chi, Quoc V. Le, and Denny Zhou. 2022. Chain-of-thought prompting elicits reasoning in large language models. In Proceedings of the 36th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’22). Curran Associates Inc., Red Hook, NY, USA, Article 1800, 14 pages.
- Widom (1996) Jennifer Widom. 1996. The starburst active database rule system. IEEE Transactions on Knowledge and Data Engineering 8, 4 (1996), 583–595.
- Willard and Louf (2023) Brandon T. Willard and Rémi Louf. 2023. Efficient Guided Generation for Large Language Models. arXiv:2307.09702 [cs.CL] https://arxiv.org/abs/2307.09702
- Xu et al. (2019) Keyulu Xu, Weihua Hu, Jure Leskovec, and Stefanie Jegelka. 2019. How Powerful are Graph Neural Networks?. In International Conference on Learning Representations. https://openreview.net/forum?id=ryGs6iA5Km
- Zheng et al. (2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric P. Xing, Hao Zhang, Joseph E. Gonzalez, and Ion Stoica. 2023. Judging LLM-as-a-judge with MT-bench and Chatbot Arena. In Proceedings of the 37th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’23). Curran Associates Inc., Red Hook, NY, USA, Article 2020, 29 pages.
- Zheng et al. (2024) Lianmin Zheng, Liangsheng Yin, Zhiqiang Xie, Chuyue Sun, Jeff Huang, Cody Hao Yu, Shiyi Cao, Christos Kozyrakis, Ion Stoica, Joseph E. Gonzalez, Clark Barrett, and Ying Sheng. 2024. SGLang: Efficient Execution of Structured Language Model Programs. arXiv:2312.07104 [cs.AI] https://arxiv.org/abs/2312.07104